Please include all details on this sheet.

Formulas must be used.

Wilkins Corporation manufactures two types of handbags—Standard and Custom—and applies manufacturing overhead to all units using machine hours.

Production information follows.

Standard Custom Anticipated volume (units) 58,500 31,200

Direct-material cost per unit $ 46 $ 69

Direct-labor cost per unit 51 60

The controller, who is studying the use of activity-based costing, has determined that the firm’s overhead can be identified with three activities: manufacturing setups, machine processing, and purchase orders.

Data on the number of setups, machine hours, and purchase orders, which are the activities’ three respective cost drivers, follow.

Standard Custom Setups 176 136

Machine hours 62,400 87,750

Purchase orders 200 150

The firm’s total overhead is: $23,423,400 It is subdivided as follows:

manufacturing setups $5,110,560

machine processing $14,054,040

purchase orders $4,258,800

Required:

1. Compute the per unit total manufacturing cost of Standard and Custom handbags by using the company’s current overhead costing procedures with machine hours.

2. Compute the per unit total manufacturing cost of Standard and Custom handbags by using activity-based costing.

3. Assume that the current selling price of a Standard handbag is $330.00 and the marketing manager is contemplating a $60 discount to stimulate volume. Is this discount advisable? Show calculations to support your answer.

4. Which method should the company use and why?
Answer #1

1. Total expected machine hours = 62,400 + 87,750 = 150,150

Predetermined overhead rate = Estimated Total Overhead / Estimated Machine Hours = $ 23, 423,400 / 150,150 = $ 156 per machine hour.

Standard Custom
Direct material $ 46 $ 69
Direct labor 51 60
Manufacturing Overhead 166.40 438.75
Unit Manufacturing Cost $ 263.40 $ 567.75

Manufacturing overhead per unit for Standard: ( 62,400 x 156 ) / 58,500 units = 166.40

Manufacturing overhead per unit for Custom : ( 87,750 x 156 ) / 31,200 units = 438.75

2.

Standard Custom
Direct material $ 46 $ 69
Direct labor 51 60
Manufacturing setup 49.28 71.40
Machine processing 99.84 263.25
Purchase orders 41.60 58.50
Unit Manufacturing Cost $ 287.72 $ 522.1

Setup costs per unit: Activity rate = $ 5,110,560 / ( 176 + 136 ) = $ 16,380 per setup.

Standard : $ 16,380 * 176 / 58,500 = 49.28

Custom : $ 16,380  * 136 / 31,200 = 71.40

3. The discount is not advisable. The supporting calculations are as follows :

Current selling price $ 330
Less: Proposed discount 60
Discounted selling price $ 270
Less: Unit Manufacturing Cost (287.72)
Gross margin ( loss ) $ ( 17.72)

4. The company should use activity based costing, as the two products are complex in terms of volume diversity, as well as diversity in their respective consumption of activities or the cost drivers. Therefore, using traditional costing ( single allocation base ) would lead to under-costing of one product at the cost of over-costing of another. Under-costing results in under-pricing, and consequent loss of profitability. Over-costing on the other hand leads to loss of market share.